OpenStudy (anonymous):

Medal for major help please :) Part 1: When writing linear equations, how do you determine which form of a line to use? Part 2: Choose 1 set of points from the choices below. Then, solve the problem and post your solution, showing your steps. Write an equation in point-slope form for the line that passes through one of the following pairs of points (you may choose the pair you want to work with). Then, use the same set of points to write the equation in standard form and again in slope-intercept form. Point pairs ◦(5, 1), (–3, 4) ◦(0, –2), (3, 2) ◦(–2, –1), (1, 2)

OpenStudy (anonymous):

:( no clue please explain and help

OpenStudy (anonymous):

@ikram002p @zaibali.qasmi @Haseeb96

OpenStudy (anonymous):

@MarrBear

OpenStudy (anonymous):

@Asevilla5

OpenStudy (anonymous):

Okay, when determiningthe line, you have to look at the Exponent, or the degree.

OpenStudy (anonymous):

select the 2nd choice...(0,-2),(3,2) so linear eq are 0x-2y=0 3x+2y=0 solving these simply addition gives x=0 y=0 and eq of line in point slope form is y=mx+c so as x=y=0 so eq reduces to c=0 and also slope intercept form gives c=0

OpenStudy (anonymous):

@zaibali.qasmi Ok so how would I answer part 1 Part 1: When writing linear equations, how do you determine which form of a line to use?

OpenStudy (anonymous):

@Asevilla5

OpenStudy (anonymous):

I don't know if this answer is correct according to a textbook, but I'd say you should use slope-intercept form when you want to graph the linear function, because it tells you the important parts of it right there.

OpenStudy (asevilla5):

yeah use the slope intercept form it's the easiest way to do it

OpenStudy (asevilla5):